OpenStudy (anonymous):

MEDAL AND FAN!!!!!! PLZ HELP ASAP!!!!! how many times larger is the new volume of a cylinder if only the height tripled, and the radius remained the same? 3 times larger 6 times larger 9times larger the volume stays the same. Thnx n advance btw.. pls tell me how u gt it..

OpenStudy (anonymous):

@igreen, @iambatman

OpenStudy (anonymous):

do you have an image of the cylinder @b.trail.coco4231?

OpenStudy (anonymous):

no,i have SOS. I am homeschooled. tht is the actual question. @AMYCARTER

OpenStudy (matt101):

The equation for the volume of a cylinder is: \[V_0= \pi r^2h\] If you triple the height, you get: \[V_{new}=\pi r^2 (3h)\] Which is the same as saying: \[V_{new}=3(\pi r^2 h)\]\[V_{new}=3V_0\] So you can see that your new volume after tripling the height is just 3 times the starting volume, V(o). Doesn't matter what the actual volume is because we just want to know how the volume changed.

OpenStudy (anonymous):

Thank u!!!!!!!!!!!! @matt101